f(x,y)=x²-xy+y²/2+3


1
Expert's answer
2021-03-31T13:48:51-0400

The question seems incomplete, however, I can guess that you are asking to find the linearization of "f(x,y)=x^2-xy+y^2\/2+3" at a given point say (3,2). If this is the case

"f(3,2)=9-6+2+3=8"

"\frac{\partial \:}{\partial \:x}\left(f\left(x,y\right)\right)=2x-y"

"\frac{\partial \:}{\partial \:y}\left(f\left(x,y\right)\right)=-x+y"

"\nabla f(x,y)=<2x-y,-x+y>"

"\nabla f(x,y)=<4,-1>"

Linearization "z-8=4(x-3)-1(y-2)"

"4x-y-12+2-z+8=0"

"4x-y-z-2=0"